# Polarimetric compressed sensing with hollow, self-assembled diffractive films

### v1

@@ 8.2 COMSOL
We calculate the transmitted fields in a hexagonal lattice with COMSOL [Fig. 14(B)i]. For incident TE and TM polarized light, we vary the azimuthal planes of incidence ($\phi=0,7.5,15,22.5,and30^{\circ}$), the polar angles of incidence ($\theta=0..60^{\circ}$), and the different shell thicknesses $t=40-120$nm. A coherent weighting coefficient for each thickness, $c(t)$, is fit to minimize mean-squared-error to the experimental TE transmission curve shown in Fig. 14(B)ii. We subsequently fit the numerically-weighted and coherently superposed TE and TM transmitted curves for this TE-fit $c(t)$, which shows good agreement with the experimental results. This indicates that the TE and TM dips in polarized transmission can, to some extent, be explained by a thin, solid shell lattice with a distribution around $t=100$-nm thickness.
@@ 8.3 Mie theory
The optical properties of homogeneous and core-shell nanospheres, such as absorption, scattering, extinction efficiencies, and scattering patterns, can be determined using Mie’s theory. The electromagnetic simulations of periodic/random arrays of core-shell spheres are a challenging task; for this reason, in some cases, it is desirable to model the core-shell structure as an equivalent sphere with an effective refractive index for simply the problem.
With Mie theory, it can be shown that individual hollow spheres exhibit 20-times reduced specular reflection and 2-3 times transmission compared to individual solid PS spheres of the same size. This would explain why the hollow-sphere geometry is important for observing the polarization-dependent transformations: a greater percentage of light is back-scattered into nonlocal modes in the hollow-sphere lattice than a PS-sphere lattice. Results from Mie theory are shown in Fig. 15
@@ 8.4 Effective medium theory
Effective medium theories provide a means to calculate the dielectric function of an equivalent sphere with the same radius and electromagnetic characteristics as a spherical core shell. Among various theories, we use the weighted average approach to compute the effective dielectric function,
@@
We simulate this 2-D photonic crystal lattice of spheres with these refractive indices with the S4 package and calculate the specular transmission and reflection of the TE (s) and TM (p) polarizations. The experimental transmission and reflective specular measurements are shown in Fig. 16 for one orientation and for an average of seven, equally-spaced rotated orientations. With one orientation, we observe what appears to be a Rayleigh anomaly at an incident angle of 15 degrees for the TE polarization. With seven averaged orientations, the feature is broadened and extends to an angle of incidence of 20 degrees. We believe this lattice effect is a key feature in the PANI-HS enabling its polarization transformations.
Figure 16: Specular reflection (top) and reflection (bottom) for (A) a single orientation and (B) average of seven rotated orientations for simulated 720-nm diameter, single-layer photonic-crystal lattice. (refractive index $n=1.26+0.025i$). The solid and dotted lines indicated simulated and experimental curves, respectively.
@@ 9 Multi-parameter estimation with shallow neural networks (SNN)
The SNN inputs are the stacked signals from $m$ detectors equally spaced 5-30 apart between $-\theta_{max}=-75^{\circ}$ and $\theta_{max}=75^{\circ}$. Each input entry to the SNN is a product of the measured $S_{2}$ and the smoothed power signal for $m$ detectors. The smoothing is provided by a Savitzky–Golay filter (SGF). The smoothing only occur over input angles since we assume that each output-angle detector is independent (that there is no signal processing coordinated between detectors except at the neural network).
The SNN model is a two-layer SNN with nonlinear “gelu” activations and loss function that is the mean absolute error of $[\theta_{i},S_{2},S_{3}]$ predictions where each parameter has a range [-1, 1]. A validation split of 0.1 is used and the variance in the prediction is calculated for 15 test splits.

### v2

@@ 8.2 COMSOL
We calculate the transmitted fields in a hexagonal lattice with COMSOL [Fig. 5(B)i]. The refractive index data from prior PANI characterization 56 is used. For incident TE and TM polarized light, we vary the azimuthal planes of incidence ($\phi=0,7.5,15,22.5$, and $30^{\circ}$), the polar angles of incidence ($\theta=0..60^{\circ}$), and the different shell thicknesses $t=40-120$nm. A coherent weighting coefficient for each thickness, $c(t)$, is fit to minimize mean-squared-error to the experimental TE transmission curve shown in Fig. 5(B)ii. We subsequently fit the numerically-weighted and coherently superposed TE and TM transmitted curves for this TE-fit $c(t)$, which shows good agreement with the experimental results. This indicates that the TE and TM dips in polarized transmission can, to some extent, be explained by a thin, solid shell lattice with a distribution around $t=100$-nm thickness.
@@ 8.3 Mie theory
The optical properties of homogeneous and core-shell nanospheres, such as absorption, scattering, extinction efficiencies, and scattering patterns, can be determined using Mie’s theory 57, 58. The electromagnetic simulations of periodic/random arrays of core-shell spheres are a challenging task; for this reason, in some cases, it is desirable to model the core-shell structure as an equivalent sphere with an effective refractive index for simply the problem.
With Mie theory, it can be shown that individual hollow spheres exhibit 20-times reduced specular reflection and 2-3 times transmission compared to individual solid PS spheres of the same size. This would explain why the hollow-sphere geometry is important for observing the polarization-dependent transformations: a greater percentage of light is back-scattered into nonlocal modes in the hollow-sphere lattice than a PS-sphere lattice. Results from Mie theory are shown in Fig. 15.
@@ 8.4 Effective medium theory
Effective medium theories provide a means to calculate the dielectric function of an equivalent sphere with the same radius and electromagnetic characteristics as a spherical core shell. Among various theories, we use the weighted average approach 59 to compute the effective dielectric function,
@@
We simulate this 2-D photonic crystal lattice of spheres with these refractive indices with the S4 package 60 and calculate the specular transmission and reflection of the TE (s) and TM (p) polarizations. The experimental transmission and reflective specular measurements are shown in Fig. 16 for one orientation and for an average of seven, equally-spaced rotated orientations. With one orientation, we observe what appears to be a Rayleigh anomaly at an incident angle of 15 degrees for the TE polarization. With seven averaged orientations, the feature is broadened and extends to an angle of incidence of 20 degrees. We believe this lattice effect is a key feature in the PANI-HS enabling its polarization transformations.
Figure 16: Specular reflection (top) and transmission (bottom) for (A) a single orientation and (B) average of seven rotated orientations for simulated 720-nm diameter, single-layer photonic-crystal lattice. (refractive index $n=1.26+0.025i$). The solid and dotted lines indicated simulated and experimental curves, respectively.
@@ 9.1 Two-layer simple model
The SNN inputs are the stacked signals from $m$ detectors equally spaced 5-30 apart between $-\theta_{max}=-75^{\circ}$ and $\theta_{max}=75^{\circ}$. Each input entry to the SNN is a product of the measured ${\rm S}_{2}$ and the smoothed power signal for $m$ detectors. The smoothing is provided by a Savitzky–Golay filter (SGF). The smoothing only occur over input angles since we assume that each output-angle detector is independent (that there is no signal processing coordinated between detectors except at the neural network).
The SNN model is a two-layer SNN with nonlinear “gelu” activations and loss function that is the mean absolute error of $[\theta_{i},{\rm S}_{2},{\rm S}_{3}]$ predictions where each parameter has a range [-1, 1]. A validation split of 0.1 is used and the variance in the prediction is calculated for 15 test splits.
9.2 Details regarding Savitzky-Golay Filtering (SGF)
@@
9.3 Definition of compression ratio
where $\Delta\theta_{o}$ is the average detector spacing, $\varepsilon_{\rm Pointing}$, $\varepsilon_{\rm S_{2}}$, and $\varepsilon_{\rm S_{3}}$ are the SNN prediction mean absolute error (MAE) associated with pointing, ${\rm S}_{2}$ and ${\rm S}_{3}$, respectively. The factors of $\pi$ in Eq. (15) refer to the area of the equatorial-plane unit circle of the Poincare sphere. Eq. (15) is an approximation, since we do not predict ${\rm S}_{0}$ and our inputs are ${\rm S}_{2}$, ${\rm S}_{3}=\pm 1$ while we consider a regression model where ${\rm S}_{2}$ and ${\rm S}_{3}$ are independent variables. For the purpose of this investigation, Eq. (15) is a reliable measure since ${C_{\rm R}}$ decreases when the polarization prediction error increases. Moreover, for a solid beam spot, the pointing MAE is approximately half of the detector spacing $\varepsilon_{\rm Pointing}\sim\Delta\theta_{o}/2$ and when $\varepsilon_{\rm S_{2}},\varepsilon_{\rm S_{3}}=0$, Eq. (15) yields ${C_{\rm R}}=2$. This factor of two is expected since the full-Stokes polarization characterization typically requires four separate measurements while our encoded approach employs only two.